OpenStudy (anonymous):

Find the oblique asymptote of f(x)= x^2+3x-4 / x+4

OpenStudy (anonymous):

Have you tried long division?

OpenStudy (anonymous):

I'm really clueless sorry >.<

OpenStudy (anonymous):

The oblique asymptote is the quotient (result) you get from long division without the remainder.  As it turns out, the numerator is perfectly factorizable: \(x^2+3x-4=(x+4)(x-1)\), so \[\frac{x^2+3x-4}{x+4}=\frac{(x+4)(x-1)}{x+4}=x-1~~~~\text{ (when }x\not=-4)\] So the oblique "asymptote" is the line \(x-1\).